### v3

# Boosting Adversarial Transferability with Low-Cost Optimization via Maximin Expected Flatness

Abstract
Transfer-based attacks craft adversarial examples on white-box surrogate models and directly deploy them against black-box target models, offering practical query-free threat scenarios. While flatness-enhanced methods have recently emerged to improve transferability by enhancing the loss surface flatness of adversarial examples, their divergent flatness definitions and heuristic attack designs suffer from unexamined optimization limitations and missing theoretical foundation, thus constraining their effectiveness and efficiency. This work exposes the severely imbalanced exploitation-exploration dynamics in flatness optimization, establishing the first theoretical foundation for flatness-based transferability and proposing a principled framework to overcome these optimization pitfalls. Specifically, we systematically unify fragmented flatness definitions across existing methods, revealing their imbalanced optimization limitations in over-exploration of sensitivity peaks or over-exploitation of local plateaus. To resolve these issues, we rigorously formalize average-case flatness and transferability gaps, proving that enhancing zeroth-order average-case flatness minimizes cross-model discrepancies. Building on this theory, we design a Maximin Expected Flatness (MEF) attack that enhances zeroth-order average-case flatness while balancing flatness exploration and exploitation. Extensive evaluations across 33 models and 43 current transfer-based attacks demonstrate MEF’s superiority: it surpasses the state-of-the-art PGN attack by 4% in attack success rate at half the computational cost and achieves 8% higher success rate under the same budget. When combined with input augmentation, MEF attains 15% additional gains against defense-equipped models, establishing new robustness benchmarks. Our code is available at https://github.com/SignedQiu/MEFAttack.
(a) VMI [1]
(b) TPA [2]
(c) FEM [3]
(d) EMI [4]
(e) GNP [5]
(f) APP [6]
(g) RAP [7]
(h) PGN [8]
(i) MEF
Figure 1: Visualization of adversarial loss landscapes for nine attacks on Res-50 [9]. The loss surfaces are constructed by perturbing adversarial examples along two random directions. Our MEF attack achieves the flattest loss landscape, indicating superior flatness-based transferability.
I Introduction
Adversarial examples [10, 11] pose a significant security threat to deep neural networks (DNNs). By introducing imperceptible perturbations to benign inputs, attackers can craft adversarial examples that mislead DNNs into producing erroneous predictions. Early adversarial attacks primarily operated under two paradigms: white-box attacks that require full access to the target model’s architecture and parameters, such as L-BFGS [10] and PGD [11], and query-based grey-box attacks that rely on iterative adversarial direction estimation through repeated model queries, as seen in Boundary Attack [12] and SimBA [13]. While effective, these methods face critical deployment barriers: white-box attacks assume unrealistic access to victim models, while query-based methods incur prohibitive query overhead, making them impractical in real-world scenarios.
Transfer-based attacks [14, 15, 16, 1, 17, 2] overcome these limitations by crafting adversarial examples on surrogate models and directly transferring them to black-box targets. This paradigm eliminates both dependency on target model’s internals and query interactions, making it a practical threat to real-world systems. Among these attacks, a recent class of flatness-enhanced methods [3, 7, 8, 2] has emerged as a promising direction by linking the geometric flatness of the adversarial loss landscape to adversarial transferability. Flatness-enhanced methods craft adversarial examples with flatter neighborhood loss landscapes, enhancing cross-model robustness and achieving state-of-the-art transferability [8].
Despite their empirical success, flatness-enhanced methods face two fundamental limitations. Firstly, they adopt fragmented flatness definitions without systematic comparison, failing to discern their inherent advantages and critical flaws. Secondly, they rely on the empirically assumed but unproven connection between flatness and transferability [7, 8, 5, 3], leading to non-rigorous and suboptimal optimization.
To address these limitations, we first unify fragmented flatness definitions within a multi-order framework that differentiates between worst-case and average-case flatness. Through experiments, we demonstrate that worst-case flatness optimization over-explores high-curvature regions and average-case methods over-exploit suboptimal plateaus, two critical flaws responsible for the performance bottleneck in existing attacks. We further establish the first theoretical proof linking flatness to transferability, formalizing that enhancing multi-order average-case flatness reduces cross-model discrepancies. Our theoretical analysis certifies zeroth-order flatness as the dominant transferability source, paving the way for high-transferability attacks with reduced computational costs.
Based on the empirical and theoretical analysis, we propose the Maximin Expected Flatness (MEF) attack to enhance adversarial transferability while reducing computational cost. Our key insight is to leverage the complementary strengths of average-case and worst-case flatness optimization, where average-case methods stabilize local exploitation while worst-case strategies preserve global exploration capability. MEF attack explicitly optimizes zeroth-order worst-neighborhood average flatness, a unified metric that internally computes neighborhood-level average-case flatness to stabilize local exploitation and externally identifies worst-case regions to guide global exploration. MEF implements this via gradient-guided neighborhood conditional sampling (NCS) for efficient worst-region identification and gradient balancing optimization (GBO) that reuses gradients across stages, reducing computation by 50% while enhancing transferability. By coupling these mechanisms, MEF achieves flattest loss landscape (as shown in Fig. 1) and enhanced adversarial transferability. Third-party evaluation on the widely-adopted TransferAttack benchmark shows MEF dominates 26 gradient-based methods as the new leader. Our main contributions are summarized as follows:
  • We unify fragmented flatness definitions across existing methods and establish the first theoretical foundation of flatness-based transferability, connecting multi-order average-case flatness with adversarial transferability.
  • We propose Maximin Expected Flatness (MEF) attack, together with Neighborhood Conditional Sampling (NCS) and Gradient Balancing Optimization (GBO) to boost adversarial transferability with efficient computation.
  • Extensive experiments and third-party evaluation [^1] Available in: https://github.com/Trustworthy-AI-Group/TransferAttack validate the superior effectiveness of MEF attack, surpassing the SOTA PGN [8] attack by 4% in attack success rate on average at half the computational cost and achieves 8% higher success rate under the same budget.
II Related Work
This section systematically reviews adversarial machine learning literature through two lenses: (1) we survey transfer-based adversarial attacks, and (2) analyze various defense mechanisms, including both reactive and proactive defenses.
II-A Adversarial Attacks
We first give the formal definition of adversarial attacks and then introduce various transfer-based attacks. Adversarial attacks craft perturbed examples $\mathbf{x}^{adv}\in B_{\epsilon}(\mathbf{x})=\{\mathbf{x}^{\prime}:\|\mathbf{x}^{\prime}-\mathbf{x}\|_{p}\leq\epsilon\}$ to fool a neural network $\mathcal{F}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathbf{x}\in\mathcal{X}$ is the clean input with label $y\in\mathcal{Y}$. The attack objective:
$\underset{\mathbf{x}^{adv}\in B_{\epsilon}(\mathbf{x})}{\max}J(\mathbf{x}^{adv},y;\mathcal{F}),$ (1)
where $J:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ denotes the adversarial loss function. Following standard practice, we adopt the $\ell_{\infty}$-norm ($p=\infty$) for perturbation constraints. For notation simplicity, we may write $J(\mathbf{x})\triangleq J(\mathbf{x},y;\mathcal{F})$. An intriguing property of adversarial examples is transferability, where adversarial examples crafted on surrogate model $\mathcal{F}$ also deceive the unknown target model $\mathcal{F}^{\prime}$, enabling black-box attacks. Transfer-based attacks can be roughly divided into data-driven methods, model-driven methods and optimization-driven methods, as follows.
Data-driven Methods. Data-driven methods [14, 15, 17, 18] boost adversarial transferability by aggregating gradients over multiple augmented variants through data transformation strategies. The key distinction between these methods centers on the specific transformation operators. DI [14] optimizes over resized images, TI [15] applies spatial shifting, and SI [17] manipulates pixel scaling. Beyond these single-input spatial transformations, Admix [18] enhances diversity by blending cross-category images, while SSA [19] perturbs frequency domains to break spatial locality. To further amplify input variety, composite strategies integrate multiple transformations [20, 21]. However, such strategies incur significant computational overheads and induce semantic distortions, ultimately compromising their effectiveness.
Model-driven Methods. Model-driven methods [22, 23, 24, 25] exploit geometric priors in surrogate models’ feature spaces to enhance adversarial transferability, primarily through two strategies: model ensemble aligns heterogeneous architectures via joint perturbation optimization to amplify spectral commonality with victims [24, 25], while surrogate refinement adversarially tunes surrogates or aligns gradients to reshape decision boundaries [22, 23]. However, model ensemble scales linearly with surrogates’ computational costs, and refinement struggles on large datasets due to complex regularization, limiting both to small-scale applications.
Optimization-driven Methods. Optimization-driven methods [16, 7, 26, 27], enhance transferability by redesigning optimization objectives [26, 27] or algorithms [28, 16]. Advanced Objective strategies replace cross-entropy loss with feature discrepancy losses that align surrogate-victim model features [26, 27]. Generative Modeling strategies leverage generative networks to synthesize perturbations without iterative updates [28]. Gradient Stabilization strategies retain simple losses but refine gradient dynamics via momentum [16, 17, 4] or variance reduction [1]. Compared to the former two paradigms requiring complex losses or architectures, gradient stabilization strategies achieve transferability with minimal complexity, avoiding auxiliary data and models.
Recent advances in gradient stabilization attacks have seen the emergence of flatness-enhanced methods, which propose that adversarial examples within flatter loss regions exhibit stronger transferability. These methods diverge into two strategies: those optimizing average-case flatness by aggregating gradients over local neighborhoods (e.g., FEM [3], APP [6]), and those pursuing worst-case flatness by minimizing gradient peaks (PGN [8]) or maximizing minimal loss (RAP [7]). However, existing works adopt fragmented flatness definitions while relying on empirically assumed correlations between flatness and transferability. Crucially, these methods typically employ flatness as a heuristic regularizer without a rigorous theoretical foundation linking specific geometric properties to the transferability gap. This theoretical void leads to suboptimal optimization directions. Our work bridges this gap by establishing unified flatness-transferability guarantees and adaptive optimization that harmonizes both flatness criteria.
II-B Adversarial Defenses
Adversarial defenses can be categorized into two paradigms: proactive defenses that strengthen model robustness through architectural or parametric adjustments, and reactive defenses that deploy external mechanisms to sanitize or detect adversarial inputs without modifying model parameters.
Proactive Defenses. Proactive defenses aim to intrinsically harden model parameters against adversarial perturbations. The dominant approach, adversarial training (AT) [11, 29], reformulates robust model optimization as a min-max optimization problem, where the inner maximization generates perturbations to craft challenging adversarial examples, while the outer minimization enforces the model to correctly classify these perturbed samples. Variants like ensemble adversarial training (EAT) [30] augment robustness through multi-model perturbations during training, strengthening defense against cross-architecture transfer attacks. In contrast to AT’s empirical robustness, randomized smoothing [31] provides certifiable robustness by injecting Gaussian noise into inputs and certifying predictions over perturbed samples. While this method theoretically guarantees robustness within a certified radius, such radii are often smaller than adversarial perturbations defended by AT, rendering it ineffective against practical adversarial threats despite its theoretical appeal.
Reactive Defenses. Reactive defenses mitigate adversarial attacks without modifying model parameters. Adversarial purification [32, 33] projects inputs onto clean data manifolds using generative models [32, 33], while adversarial detection identifies adversarial examples via activation anomalies or confidence thresholds. These strategies offer plug-and-play deployment but suffer from two notable limitations: purification efficacy depends critically on generative model capacity [34, 35], and detectors can be bypassed by sophisticated adaptive attacks that mimic clean data statistics.
III Limitations in Flatness Enhancement
The heuristic adoption of flatness concepts in flatness-enhanced methods lacks rigorous theoretical grounding while suffering from absence of systematic comparisons between flatness metrics, which constrains their effectiveness and efficiency. In this section, we first formalize flatness as a measurable geometric property, resolving definitional ambiguities across existing methods. Then we identify two inherent limitations in current flatness-enhanced methods. Finally, we distill the critical yet understudied challenges in balancing local exploitation and global exploration for real-world deployment.
III-A K-th Order Flatness
TABLE I: Comparative analysis of flatness-enhanced methods, summarizing their conference venues, optimization target, and flatness formulations. We systematically compare six representative methods’ mathematical frameworks for achieving flatness.
Flatness-Enhanced Method Conference (Year) Optimization Target Flatness Formulation
Reverse Adversarial Perturbation (RAP [7]) NeurIPS 2022 $\underset{\mathbf{x}^{adv}\in{B}_{\epsilon}(\mathbf{x})}{max}\,\underset{\mathbf{x^{\prime}}\in{B}_{\xi}(\mathbf{x}^{adv})}{min}\,J(\mathbf{x^{\prime}})$ $\widehat{R}^{(0)}_{\xi}(\mathbf{x})$
Gradient Norm Penalty (GNP [5]) ICIP 2022 $\underset{\mathbf{x}^{adv}\in{B}_{\epsilon}(\mathbf{x})}{max}\,\underset{\mathbf{x^{\prime}}\in{B}_{\xi}(\mathbf{x}^{adv})}{min}\,[J(\mathbf{x}^{adv})-\lambda\cdot{\|\nabla_{\mathbf{x^{\prime}}}J(\mathbf{x^{\prime}})\|_{2}}]$ $\widehat{R}^{(1)}_{\xi}(\mathbf{x})$
Penalizing Gradient Norm (PGN [8]) NeurIPS 2023 $\underset{\mathbf{x}^{adv}\in{B}_{\epsilon}(\mathbf{x})}{max}\,\underset{\mathbf{x^{\prime}}\in{B}_{\xi}(\mathbf{x}^{adv})}{min}\,[J(\mathbf{x}^{adv})-\lambda\cdot{\|\nabla_{\mathbf{x^{\prime}}}J(\mathbf{x^{\prime}})\|_{2}}]$ $\widehat{R}^{(1)}_{\xi}(\mathbf{x})$
Flatness Enhanced Method (FEM [3]) ICME 2023 $\underset{\mathbf{x}^{adv}\in{B}_{\epsilon}(\mathbf{x})}{max}\,\mathbb{E}_{\mathbf{x^{\prime}}\sim\mathrm{Unif}{B}_{\xi}(\mathbf{x}^{adv})}\,\left[J(\mathbf{x^{\prime}})\right]$ $\overline{R}^{(0)}_{\xi}(\mathbf{x})$
Adversarial Pixel Perturbation (APP [6]) ICIP 2023 $\underset{\mathbf{x}^{adv}\in{B}_{\epsilon}(\mathbf{x})}{max}\,\mathbb{E}_{\mathbf{x^{\prime}}\sim\mathrm{Unif}{B}_{\xi}(\mathbf{x}^{adv})}\,\left[J(\mathbf{x^{\prime}})\right]$ $\overline{R}^{(0)}_{\xi}(\mathbf{x})$
Theoretically Provable Attack (TPA [2]) NeurIPS 2024 $\underset{\mathbf{x}^{adv}\in{B}_{\epsilon}(\mathbf{x})}{max}\,\mathbb{E}_{\mathbf{x^{\prime}}\in{B}_{\xi}(\mathbf{x}^{adv})}\,{\left[\|\nabla_{\mathbf{x^{\prime}}}J(\mathbf{x^{\prime}})\|_{2}\right]}$ $\overline{R}^{(1)}_{\xi}(\mathbf{x})$
(a) RAP [7] ($\widehat{R}^{(0)}_{\xi}(\mathbf{x})$)
(b) FEM [3] ($\overline{R}^{(0)}_{\xi}(\mathbf{x})$)
(c) PGN [8] ($\widehat{R}^{(1)}_{\xi}(\mathbf{x})$)
(d) TPA [2] ($\overline{R}^{(1)}_{\xi}(\mathbf{x})$)
[…]
In this section, we introduce the formal definitions of k-order worst- and average-case flatness metrics, establishing a unified taxonomy for flatness enhancement and enabling systematic classification of six flatness-enhanced methods.
The study of loss landscape flatness originated in model generalization theory, where flatter minima were shown to correlate with improved generalization performance [38, 39, 40]. In model generalization, the most popular mathematical definition of flatness, proposed by SAM [38], considers the maximal loss value within a radius, termed the zeroth-order worst-case flatness. Building upon this foundation, Gradient norm Aware Minimization (GAM) [40] extends this framework to first-order worst-case flatness, a stricter geometric criterion that constrains maximal gradient norms rather than loss values.
Based on SAM’s framework [38] for minimizing sharpness, RAP [7] adapts the zeroth-order worst-case flatness to construct transferable adversarial examples. Motivated by RAP [7], Penalizing Gradient Norm (PGN) [8] enhances adversarial transferability by enhancing the first-order worst-case flatness. TPA [2] further correlates transferability with first-order average flatness (local mean gradient norms) and second-order gradient component. We give the definition of $\xi$-radius $n$-order worst-case flatness $\widehat{R}^{(n)}_{\xi}(\mathbf{x})$ and average-case flatness $\overline{R}^{(n)}_{\xi}(\mathbf{x})$ in Definition 1. Unless otherwise specified, the zeroth‑order derivative is taken to be the loss function itself, i.e. $\nabla^{0}J(\mathbf{x})=J(\mathbf{x})$.
Definition 1 ($\xi$-radius $n$-order Flatness).
For any $\xi>0$ and integer $k\in\mathbb{N}$, the worst- and average-case $\xi$-radius $n$-order flatness of the loss $J(\mathbf{x})$ at sample $\mathbf{x}$ are defined as follows:
  • Worst-case flatness:
    $\widehat{R}^{(n)}_{\xi}(\mathbf{x})\triangleq\max_{\mathbf{x}^{\prime}\in B_{\xi}(\mathbf{x})}\big\|\nabla^{n}J(\mathbf{x}^{\prime})-\nabla^{n}J(\mathbf{x})\big\|_{2},$ (2)
  • Average-case flatness:
    $\overline{R}^{(n)}_{\xi}(\mathbf{x})\triangleq\mathbb{E}_{\mathbf{x}^{\prime}\sim\mathrm{Unif}(B_{\xi}(\mathbf{x}))}\big[\big\|\nabla^{n}J(\mathbf{x}^{\prime})-\nabla^{n}J(\mathbf{x})\big\|_{2}\big],$ (3)
where $\mathbf{x}^{\prime}$ denotes a point in $\ell_{\infty}$-ball $B_{\xi}(\mathbf{x})=\big\{\mathbf{x}^{\prime}\in\mathbb{R}^{d}\,\big|\,\|\mathbf{x}^{\prime}-\mathbf{x}\|_{\infty}\leq\xi\big\}$ centered at $\mathbf{x}$ with radius $\xi$, $\nabla^{n}J(\mathbf{x})$ denotes the nn‑th order derivative of $J$ at $\mathbf{x}$ and $\mathrm{Unif}(B_{\xi}(\mathbf{x}))$ denotes the uniform distribution on the set $B_{\xi}(\mathbf{x})$.
[…]
To our knowledge, this is the first work that unifies flatness definitions across flatness-enhanced methods and provides a taxonomy distinguishing worst/average-case flatness criteria. While prior work [41] quantifies flatness in parameter space for model robustness, our focus centers on sample-wise loss landscape flatness, a distinct perspective aligned with adversarial transferability rather than robust generalization.
III-B Rethinking Flatness-Enhanced Methods
Despite the prevalent use of distinct flatness criteria, existing works neither analyze their geometric discrepancies nor characterize their inherent limitations, leading to obscured distinctions, limited performance and suboptimal efficiency. In this section, we analyze the optimization dynamics of different flatness formulation and validate the unrevealed over-exploitation and over-exploration issues in current methods.
[…]
To validate the over-exploitation and over-exploration issues, we generate adversarial examples on Res-50 [9] using four flatness-enhanced methods (RAP [7], FEM [3], PGN [8], TPA [2]) for 100 iterations. We measure inter-update gradient cosine similarity during optimization and evaluate transfer attack success rates against Inc-v3 [37]. The gradient similarity serves as an indicator of the optimization dynamics: high similarity signifies consistent update directions for steady exploitation, whereas low values reflect frequent shifts in search directions for broad exploration.
The experimental results, illustrated in Fig. 2, reveal a clear empirical divergence. Worst-case flatness optimization exhibit significantly lower inter-iteration gradient similarities (mean similarity: RAP 0.10, PGN 0.25) compared to average-case counterparts (FEM 0.38, TPA 0.88) after convergence. The persistently low gradient similarity in worst-case optimization reveals an unstable oscillatory trajectory, an indicator of over-exploration driven by the hypersensitivity of the maximization objective to shifting local extrema in non-convex landscapes. In contrast, high similarity in average-case optimization suggests over-exploitation of local gradient consensus, leading to premature convergence on suboptimal plateaus as the expectation operator traps the search in dominant local directions. Ultimately, both pitfalls lead to premature stagnation of optimization, thereby strictly constraining the maximum achievable adversarial transferability.
III-C Design Challenges
Existing flatness-enhanced methods face two unresolved challenges: theoretical ambiguity and optimization imbalance. First, the heuristic adoption of flatness metrics lacks rigorous theoretical grounding to connect geometric flatness with adversarial transferability [7, 8], with some work even questioning their inherent correlation [2], leading to inconsistent metric selection and suboptimal performance. Second, as shown in Section III-B, current methods suffer from a critical exploitation-exploration trade-off: average-case flatness optimization over-exploits local gradient consensus, causing premature convergence to suboptimal plateaus, while worst-case criteria induce over-exploration of non-informative sensitivity peaks, resulting in oscillatory updates. Addressing these challenges requires developing a unified framework that integrates theoretical insights into flatness-transferability relationships with adaptive mechanisms to harmonize local exploitation and global exploration, thereby maximizing adversarial transferability.
IV Theoretical Analysis
Having formalized flatness metrics (Section III), we now resolve the first key challenge: the ambiguous relationship between geometric flatness and adversarial transferability. We first formally define transferability via the adversarial loss gap, then prove that improved multi-order flatness enhances it. This theory not only addresses the heuristic limitations of prior work but also provides principled optimization criteria.
IV-A Adversarial Transferability Formulation
To formally characterize transferability, we propose the Adversarial Transferability Gap (ATG). We premise this definition on three explicit assumptions regarding the surrogate $\mathcal{F}$ and target $\mathcal{F^{\prime}}$: (1) they share the same label space; (2) they employ the same loss function $J(\cdot)$; and (3) they exhibit aligned performance on clean samples (i.e., $J(\mathbf{x},y;\mathcal{F})\approx J(\mathbf{x},y;\mathcal{F^{\prime}})$).
Definition 2 (Adversarial Transferability Gap).
Given the aligned models $\mathcal{F}$ and $\mathcal{F^{\prime}}$ (as assumed above), the ATG for an input–label pair $(\mathbf{x},y)$ and perturbation $\boldsymbol{\delta}$ is defined as:
$\operatorname{ATG}\bigl((\mathbf{x},y),\boldsymbol{\delta};\mathcal{F},\mathcal{F^{\prime}}\bigr)\triangleq J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F^{\prime}})-J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F}).$ (4)
Remark A detailed justification of this metric’s reasonableness and a discussion of its limitations under assumption violations are provided in Section -A of the Supplementary Material.
[…]
IV-B Flatness and Transferability
Based on the definition established in Section IV-A, particularly in the typical scenario where transferability degrades (i.e., $ATG<0$), achieving successful adversarial transfer requires simultaneously addressing two critical objectives: (1) maximizing the adversarial loss $J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F})$ on the surrogate model $\mathcal{F}$ to strengthen the attack,, and (2) minimizing $|\operatorname{ATG}|$ to align adversarial effects between the surrogate $\mathcal{F}$ and the target $\mathcal{F^{\prime}}$. While gradient-ascent attacks (e.g., PGD [43]) readily achieve the first, systematically reducing $|\operatorname{ATG}|$ remains theoretically understudied and challenging. We therefore make this our theoretical focus.
The minimization of $|\operatorname{ATG}|$ hinges on characterizing the geometric alignment of adversarial perturbations between divergent loss landscapes. We address this by establishing a multi-order flatness-dependent bound on $|\operatorname{ATG}|$ through Theorem 1, which directly connects transferability gaps to three measurable geometric properties: surrogate model flatness $\bar{R}^{(n)}_{\xi}(\mathbf{x};F)$, Target model flatness $\bar{R}^{(n)}_{\xi}(\mathbf{x};F^{\prime})$, and Cross-model gradient discrepancy $C_{n}(\mathbf{x})$.
Theorem 1 (Flatness-based Bound on Transferability).
Let $\mathcal{F}$ and $\mathcal{F^{\prime}}$ be two models with loss $J(\mathbf{x},y;\,\cdot\,)$. Assume $J$ is locally approximable by a Taylor expansion of order $N$. For $0\leq n\leq N$, define
$C_{n}(\mathbf{x})=\mathbb{E}_{\mathbf{x^{\prime}}\sim\mathrm{Unif}(B_{\xi}(\mathbf{x}))}\|\nabla^{n}J(\mathbf{x^{\prime}},y;F^{\prime})-\nabla^{n}J(\mathbf{x^{\prime}},y;F)\|,$ (5)
and let $\bar{R}^{(n)}_{\xi}(\mathbf{x};F)$ and $\bar{R}^{(n)}_{\xi}(\mathbf{x};F^{\prime})$ be the $n$-order flatness for $F$ and $F^{\prime}$. Then for perturbation $\|\boldsymbol{\delta}\|\leq\xi$, the adversarial transferability gap satisfies
$\begin{split}\bigl|\mathrm{ATG}((\mathbf{x},y),\boldsymbol{\delta};\mathcal{F},F^{\prime})\bigr|&\leq\sum_{n=0}^{N}\frac{\|\boldsymbol{\delta}\|^{n}}{n!}\Bigl[\bar{R}^{(n)}_{\xi}(\mathbf{x};F)\\ &\qquad+C_{n}(\mathbf{x})+\bar{R}^{(n)}_{\xi}(\mathbf{x};F^{\prime})\Bigr]\\ &\qquad+\Delta R_{N}(\mathbf{x},\boldsymbol{\delta}),\end{split}$ (6)
where $\Delta R_{N}$ bounds the difference in higher-order residuals.
The proof of Theorem 1 (see Section -B in Supplementary Material) employs finite-order multivariate Taylor expansions to decompose $\operatorname{ATG}$ into curvature alignment terms plus a residual bound. This framework reveals that suppressing multi-order flatness variations and gradient discrepancies ($C_{n}$) minimizes the transferability gap. Crucially, $C_{n}(\mathbf{x})$ quantifies intrinsic surrogate-target dissimilarity; a larger $C_{n}$ loosens the bound, acting as a natural barrier to transfer attack success.
In transfer-based attacks where only surrogate $\mathcal{F}$ is accessible, optimizing source flatness $\bar{R}^{(n)}_{\xi}(\mathbf{x};F)$ becomes the sole effective strategy, as target flatness and $C_{n}(\mathbf{x})$ are uncontrollable. While TPA [2] linked flatness to transferability, it was limited to first-order analysis ($n=1$). Our theory incorporates higher-order interactions ($n\geq 1$), formalizing why heuristics [7, 5, 8, 3, 6] succeed. Theorem 1 also shows that high-order terms diminish rapidly due to the $\frac{\|\boldsymbol{\delta}\|^{n}}{n!}$ coefficient. For ReLU networks where high-order derivatives vanish ($N\approx 1$), the zeroth-order flatness $\bar{R}^{(0)}_{\xi}(\mathbf{x};F)$ becomes the dominant factor, establishing it as an efficient optimization target.
Theorem 1 also demonstrates that the impact of higher-order terms ($n\geq 1$) diminishes exponentially due to the $\frac{\|\boldsymbol{\delta}\|^{n}}{n!}$ coefficient under $\ell_{p}$-bounded perturbations ($\|\boldsymbol{\delta}\|\ll 1$), rendering zeroth-order flatness $\bar{R}^{(0)}_{\xi}(\mathbf{x};F)$, which governs expected loss variations in $B_{\xi}(\mathbf{x})$, the dominant contributor to transferability. This establishes zeroth-order flatness optimization as an efficient mechanism for maximizing adversarial transferability with theoretical guarantees.
V Method
(a) The Workflow of MEF Optimization
(b) Geometric Interpretation
Figure 3: Overview of the proposed MEF attack. (a) Illustration of the iterative optimization process. At step t1t-1, Neighborhood Conditional Sampling (NCS) transforms random samples ($x^{r}$) into worst-case neighborhood samples ($x^{n}$) via inner minimization. Outer maximization then aggregates gradients from these regions to update the adversarial example $x_{t}^{adv}$. (b) Geometric interpretation showing how MEF seeks flat maxima to minimize the Adversarial Transferability Gap (ATG) between the source model $F$ (solid line) and the target model $F^{\prime}$ (dashed line), as opposed to sharp maxima which lead to larger gaps.
Building on the theoretical linkage between multi-order flatness and transferability established in Section IV, we now address the second challenge outlined in Section III-C: designing a principled mechanism to balance local exploitation and global exploration. Instead of heuristically adopting existing metrics, our framework is strictly guided by Theorem 1: since the impact of higher-order terms ($n\geq 1$) diminishes exponentially due to the $\frac{\|\delta\|^{n}}{n!}$ coefficient, we prioritize the theoretically dominant zeroth-order flatness. Crucially, to resolve the exploitation-exploration dilemma inherent in standard definitions, we define a novel Zeroth-order Worst-Neighborhood Average Flatness and optimize it through maximin expected flatness, neighborhood conditional sampling, and gradient balancing optimization.
V-A Maximin Expected Flatness
Despite the computational efficiency and desirable theoretical properties of zeroth-order flatness metrics, existing zeroth-order flatness metrics exhibit polarized limitations: worst-case flatness $\widehat{R}^{(0)}_{\xi}(\mathbf{x})$ risks over-exploration by emphasizing local peaks, while average-case flatness $\overline{R}^{(0)}_{\xi}(\mathbf{x})$ tends toward over-exploitation by averaging out critical geometric features. To resolve the conflicting demands of local exploitation and global exploration in flatness optimization, we propose a novel zeroth-order worst-neighborhood average flatness metric that strategically integrates the stability of $\overline{R}^{(0)}_{\xi}(\mathbf{x})$ with the exploratory power of $\widehat{R}^{(0)}_{\xi}(\mathbf{x})$.
Our key insight lies in a hierarchical decomposition of flatness measurement that first identifies high-uneven sub-regions within an expanded neighborhood, then applies average-case smoothing to stabilize gradient estimation. Formally, we define the zeroth-order worst-neighborhood average flatness as:
$\widehat{\overline{R}}^{(0)}_{\gamma,\xi}(\mathbf{x})\triangleq\max_{\mathbf{x}^{\rm c}\in B_{\gamma}(\mathbf{x})}\overline{R}^{(0)}_{\xi}(\mathbf{x}^{\rm c}),$ (7)
where $\gamma$ controls the exploration radius for identifying critical subregion centers $\mathbf{x}^{\rm c}$, and $\xi$ governs the neighborhood radius for zeroth-order average-case flatness computation. This hierarchical structure explicitly separates exploration scope ($\gamma$) from local smoothness ($\xi$), where larger $\gamma$ values permit discovery of distant high-loss regions while smaller $\xi$ values ensure stable gradient estimation within identified sub-regions.
With the proposed $\widehat{\overline{R}}^{(0)}_{\gamma,\xi}$, we establish the adversarial example generation framework guided by Theorem 1. As analyzed in Section IV-A, successful transfer attacks require: (i) minimizing $|\operatorname{ATG}|$ to align adversarial effects across models, and (ii) maximizing $J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F})$ to elevate the upper bound of target model loss $J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F^{\prime}})$. To unify these objectives, we formulate the attack generation as a multi-objective optimization problem:
[…]
Substituting Definition 7 into (8) yields:
$\min_{\mathbf{x}^{\rm adv}\in B_{\epsilon}(\mathbf{x})}\Bigl\{$ $\max_{\mathbf{x}^{\rm c}\in B_{\gamma}(\mathbf{x}^{\rm adv})}\mathbb{E}_{\mathbf{x}^{\prime}\sim\mathrm{Unif}(B_{\xi}(\mathbf{x}^{\rm c}))}\bigl[\bigl|J(\mathbf{x}^{\prime})-J(\mathbf{x}^{\rm c})\bigr|\bigr]$
$-J(\mathbf{x}^{\rm adv})\Bigr\}.$ (9)
Under the $\ell_{\infty}$-norm perturbation constraint $\|\mathbf{x}^{\prime}-\mathbf{x}^{\rm c}\|_{\infty}\leq\xi$, $\|\mathbf{x}^{\rm adv}-\mathbf{x}^{\rm c}\|_{\infty}\leq\gamma$, we observe:
$J(\mathbf{x}^{\rm adv})-J(\mathbf{x}^{\rm c})\approx 0,J(\mathbf{x}^{\prime})-J(\mathbf{x}^{\rm c})\leq 0.$ (10)
Substituting (10) into (V-A), we derive the final optimization objective:
$\max_{\mathbf{x}^{\rm adv}\in B_{\epsilon}(\mathbf{x})}\min_{\mathbf{x}^{\rm c}\in B_{\gamma}(\mathbf{x}^{\rm adv})}\mathbb{E}_{\mathbf{x}^{\prime}\sim\mathrm{Unif}(B_{\xi}(\mathbf{x}^{\rm c}))}J(\mathbf{x}^{\prime}),$ (11)
where the expectation term $\mathbb{E}\bigl[J(\mathbf{x}^{\prime})\bigr]$ represents the expected flatness over local neighborhoods. This maximin structure motivates our method’s name: Maximin Expected Flatness (MEF), which strategically explores worst-case neighborhoods while exploiting average loss landscapes.
V-B Neighborhood Conditional Sampling
Directly solving the inner minimization in (11) is computationally infeasible due to the continuous loss landscape and infinite candidates in $B_{\gamma}(\mathbf{x}^{\rm adv})$. To efficiently approximate this intractable problem and obtain representative sampling points capturing worst-case flatness, we propose Neighborhood Conditional Sampling (NCS) method that leverages gradients for guided exploration.
The NCS uniformly samples $N$ points $\{\mathbf{x}_{0,i}\}_{i=1}^{N}$ within $B_{\xi}(\mathbf{x})$. At each iteration $t$, gradients $\nabla_{\mathbf{x}}J(\mathbf{x}_{t,i},y;\mathcal{F})$ are computed for all points, followed by a sign-based update:
$\mathbf{x}_{t+1,i}=\mathbf{x}_{t,i}-\xi\cdot\mathrm{sign}\left(\nabla_{\mathbf{x}}J(\mathbf{x}_{t,i},y;\mathcal{F})\right),$ (12)
which prioritizes directional consistency over gradient magnitudes. Updated points are then projected back to $B_{\xi}(\mathbf{x}_{0,i})$ and $B_{\gamma}(\mathbf{x})$ to enforce spatial constraints. After $T$ iterations, the final set $\{\mathbf{x}_{T,i}\}_{i=1}^{N}$ captures regions with minimal expected loss. This approach eliminates sampling overhead by leveraging gradients, reducing computation while maintaining efficiency.
V-C Gradient Balancing Optimization
[…]
The nested maximin structure in (11) presents an optimization imbalance: the inner minimization requires fine-grained gradient descent across $N$ sampling points, while the outer maximization must update a single adversarial example $\mathbf{x}^{\rm adv}$ to elevate losses globally. This asymmetry arises because precise inner updates demand multiple iterations, whereas the outer optimization struggles to coordinate perturbations that simultaneously affect all minimized sub-regions. Experiments with gradient descent (1–4 inner steps) on 100 ImageNet images (Res-50→Inc-v3) confirm the issue: stricter inner optimization increases required convergence steps while reducing transfer attack success rates (Figure 4).
To address the imbalanced maximin optimization problem, we propose Gradient Balancing Optimization (GBO), which strategically weakens the inner minimization’s dominance by reusing outer maximization gradients. Let $\{g_{t,i}\}_{i=1}^{N}=\nabla_{\mathbf{x}}J({\{\mathbf{x^{\prime}}_{t,i}\}}_{i=1}^{N},y;\mathcal{F})$ denote the gradients from the outer maximization at iteration $t$. Instead of computing separate gradients for inner minimization, GBO approximates them as:
$\nabla_{\mathbf{x}_{t,i}^{\rm{}^{\prime}}}J(\mathbf{x}_{t,i}^{\rm{}^{\prime}},y;\mathcal{F})\approx g_{t-1,i},$ (13)
leveraging the geometric proximity between consecutive adversarial examples. This shared gradient scheme entirely eliminates the computational overhead of inner minimization by reusing outer gradients for inner updates, accelerates convergence through aligned optimization directions, and improves transfer attack success rates.
To stabilize gradient directions, we further introduce momentum:
$\{g^{inner}_{t,i}\}_{i=1}^{N}=\frac{\{g_{t,i}\}_{i=1}^{N}}{\lvert\lvert\{g_{t,i}\}_{i=1}^{N}\rvert\rvert}_{1}-\mu_{inner}\cdot\{g^{inner}_{t,i}\}_{i=1}^{N},$ (14)
where $\mu_{inner}$ controls historical gradient contributions. We compare GBO_M (momentum-enhanced) and GBO_V (vanilla) against PGD-based inner minimization with different iteration steps (1–4) on 100 ImageNet images. As shown in Fig. 4, GBO achieves faster convergence and higher transfer success rates than gradient-intensive PGD baselines, with GBO_M (momentum-enhanced) outperforming GBO_V (vanilla) in attack success rates. This demonstrates that gradient reuse with momentum stabilization not only resolves the optimization imbalance but also maximizes transferability, making GBO_M our default solution.
Algorithm 1 Maximin Expected Flatness (MEF) Attack
1:Surrogate model $\mathcal{F}$ and the loss function $J$; Raw example x with ground-truth label $y$; The perturbation magnitude $\epsilon$; the maximum iterations $T$; the outer/inner decay factor $\mu_{outer}$$\mu_{inner}$; the number of sampled examples, NN; neighborhood radius $\xi$, exploration radius $\gamma$ and $U_{\gamma}(\cdot)$ is the uniform distribution over the $\ell_{\infty}$-ball $B_{\gamma}(\cdot)$.
2:An adversarial example $\mathbf{x}^{adv}$
[…]
4:for $t=0,1,\ldots,T-1$ do
5:  # Neighborhood Conditional Sampling
6:  ${\{\mathbf{x}_{t,i}\}}_{i=1}^{N}\sim U_{\gamma}(\mathbf{x}^{adv}_{t})$
7:  ${\{\mathbf{x^{\prime}}_{t,i}\}}_{i=1}^{N}={\{\mathbf{x}_{t,i}\}}_{i=1}^{N}+\xi\cdot sign(\{g^{inner}_{t,i}\}_{i=1}^{N})$
8:  # Gradient Calculation
9:  $\{g_{t,i}\}_{i=1}^{N}=\nabla_{\mathbf{x}}J({\{\mathbf{x^{\prime}}_{t,i}\}}_{i=1}^{N},y;\mathcal{F})$
10:  # Gradient Balancing Optimization
11:  $\{g^{inner}_{t,i}\}_{i=1}^{N}=\frac{\{g_{t,i}\}_{i=1}^{N}}{\lvert\lvert\{g_{t,i}\}_{i=1}^{N}\rvert\rvert}_{1}-\mu_{inner}\cdot\{g^{inner}_{t,i}\}_{i=1}^{N}$
12:  # Outer Gradient Update
13:  $g^{outer}_{t}=\mu_{outer}\cdot\ g^{outer}_{t}+\frac{1}{N}\sum_{i=1}^{N}{\frac{\{g_{t,i}\}_{i=1}^{N}}{\lvert\lvert\{g_{t,i}\}_{i=1}^{N}\rvert\rvert}_{1}}$
14:  # Adversarial Example Update
15:  $\mathbf{x}^{adv}_{t+1}=\Pi_{{B}_{\epsilon}(\mathbf{x})}[\mathbf{x}^{adv}_{t}+\alpha\cdot sign(g_{t}^{outer})]$;
16:end for
17:return $\mathbf{x}^{adv}_{T}$.
V-D Algorithm
The proposed framework integrates the objective of Maximin Expected Flatness, Neighborhood Conditional Sampling, and Gradient Balancing Optimization through a unified max-min bi-level algorithm summarized in Algorithm 1. To clarify the interaction between these components, the overall workflow is visualized in Fig. 3. By strategically sharing normalized gradients between inner minimization and outer maximization, the algorithm achieves an optimal balance between computational efficiency and attack effectiveness.
VI Experiments
TABLE II: Transfer attack success rate ($\%$) comparison of gradient-stabilized attacks on normally trained models. The best results are bold and the second best are underlined.
Attack Res-50 $\Longrightarrow$ Res-101 $\Longrightarrow$
Res-101 Inc-v3 Inc-v4 IncRes-v2 VGG-19 Dense-121 Xcept AVG Res-50 Inc-v3 Inc-v4 IncRes-v2 VGG-19 Dense-121 Xcept AVG
EMI [4] 99.3 75.0 71.2 54.3 94.8 97.2 74.3 80.9 99.8 76.8 72.6 58.5 92.9 97.2 74.0 81.7
APP [6] 99.3 83.1 79.6 73.1 95.3 97.8 83.3 87.4 99.4 83.2 80.9 72.8 93.4 96.6 82.0 86.9
FEM [3] 99.8 89.4 86.3 78.8 97.4 98.7 87.7 91.2 99.8 89.1 86.4 77.6 95.4 98.0 87.3 90.5
PGN [8] 99.2 86.3 83.9 79.0 95.6 97.8 86.9 89.8 99.5 86.8 84.5 80.0 94.5 97.1 88.1 90.1
ANDA [44] 99.4 81.5 80.1 74.1 90.7 94.7 82.2 86.1 99.5 82.6 81.2 76.4 91.0 92.8 83.7 86.7
FGSRA [45] 98.4 85.3 83.2 78.0 94.6 97.3 86.0 89.0 98.6 86.0 83.9 79.4 93.9 96.3 87.3 89.3
GI [46] 99.5 82.4 80.5 75.1 91.6 94.8 83.1 86.7 99.4 83.2 81.4 76.7 91.2 93.5 84.4 87.1
MUMODIG [47] 99.3 84.4 82.5 77.1 93.6 96.8 85.1 88.4 99.4 85.2 83.4 78.7 93.2 95.5 86.4 88.8
GAA [48] 98.5 82.5 81.1 75.1 91.7 95.7 83.2 86.8 98.9 83.6 82.2 77.4 92.0 93.8 84.7 87.5
FoolMix [49] 99.1 80.4 78.5 73.1 89.6 92.8 81.1 84.9 99.4 81.2 79.4 74.7 89.2 91.5 82.4 85.4
MEFH 99.7 90.6 88.7 82.7 98.3 99.4 90.4 92.8 99.7 91.2 88.2 82.2 97.0 98.4 90.0 92.4
MEFF 99.9 95.1 94.2 91.3 99.2 99.8 94.7 96.3 99.9 95.8 95.2 91.8 98.9 99.4 95.7 96.7
Attack Inc-v3 $\Longrightarrow$ Inc-v4 $\Longrightarrow$
Res-50 Res-101 Inc-v4 IncRes-v2 VGG-19 Dense-121 Xcept AVG Res-50 Res-101 Inc-v3 IncRes-v2 VGG-19 Dense-121 Xcept AVG
EMI [4] 78.8 74.3 78.5 72.2 79.5 80.5 78.6 77.5 79.0 75.3 86.5 71.9 85.4 82.1 82.7 80.4
APP [6] 73.5 68.0 79.2 73.1 74.7 75.7 78.5 74.7 69.0 64.7 77.1 73.5 74.7 72.5 80.7 73.2
FEM [3] 82.1 77.1 87.2 82.1 81.8 82.1 84.5 82.4 78.3 73.9 86.0 78.9 81.6 80.4 86.9 80.9
PGN [8] 84.6 79.9 88.4 84.9 84.6 86.9 89.1 85.5 84.9 82.3 91.1 85.8 89.0 87.6 90.8 87.4
ANDA [44] 81.1 76.6 85.2 80.6 80.8 81.9 85.2 81.6 80.9 77.7 86.3 81.0 84.2 83.8 87.5 83.1
FGSRA [45] 84.0 79.3 87.8 84.1 83.9 85.9 88.4 84.8 84.1 81.4 90.1 84.8 88.0 86.9 90.2 86.5
GI [46] 81.3 76.7 85.3 81.3 81.2 82.9 85.7 82.1 81.4 78.5 87.2 81.9 85.1 84.2 87.6 83.7
MUMODIG [47] 83.3 78.7 87.3 83.3 83.2 84.9 87.7 84.1 83.4 80.5 89.2 83.9 87.1 86.2 89.6 85.7
GAA [48] 82.1 77.6 86.2 81.6 81.8 82.9 86.2 82.6 81.9 78.7 87.3 82.0 85.2 84.8 88.5 84.1
FoolMix [49] 79.3 74.7 83.3 79.3 79.2 80.9 83.7 80.1 79.4 76.5 85.2 79.9 83.1 82.2 85.6 81.7
MEFH 85.8 80.8 89.4 86.1 85.2 87.4 89.9 86.4 86.2 82.4 92.8 86.1 90.1 88.5 90.8 88.1
MEFF 92.6 91.5 96.5 94.3 92.8 93.4 95.8 93.8 91.4 90.3 94.5 91.2 92.2 91.9 95.1 92.4
VI-A Experimental Setup
Dataset. We conduct experiments on the ImageNet-compatible dataset consisting of 1,000 images (299×299 resolution with annotated labels), which serves as the standard benchmark in previous works[50, 51, 24, 52, 7, 8].
Models. We select Inception-v3 (Inc-v3) [37], Inception-v4 (Inc-v4) [53], ResNet-50 (Res-50) [9] and ResNet-101 (Res-101) [9] with TorchVision-pretrained weights [54] for generating adversarial examples. Transferability evaluation covers: (1) Standard CNNs: Inc-v3 [37], Inc-v4 [53], Res-50 [9], Res-101 [9], Inception-ResNetv2 (IncRes-v2) [53], DenseNet-121 (Dense-121) [55], VGG-19bn (VGG-19) [56] and Xception (xcept) [57]; (2) Cross-architectures: MobileNet-v2 (MobileNet) [58], PNASNet-5-Large (PNASNet-L) [59], ViT-Base/16 (ViT-B/16) [60], ViT-Large/32 (ViT-L/32) [60], PiT-S [61], MLP-Mixer [62] and ResMLP [63]; (3) Defended models: Adversarially trained Inc-v3adv [29], Inc-v3ens3 [30], Inc-v3ens4 [30] and IncRes-v2ens [30]; adversarial purification defenses HGD [32] and NRP [33]; certifiable defense RS [31]; diffusion-based purifier DiffPure [34] and DensePure [35] and state-of-the-art adversarially robust models from RobustBench [64]: MIMIR [65], ARES [66] and MeanSparse [67]. (4) Multimodal Architectures (Vision-Language Models): To investigate adversarial transferability towards emerging multimodal paradigms, we evaluate VLMs including LLaVA [68], Unidiffuser [69], MiniGPT-4 [70] and Qwen2.5-VL [71].
Compared Methods. We evaluate against 43 transfer-based adversarial attacks spanning three categories: (1) 11 data-driven methods covering input augmentation (DI [14], TI [15], SI [17], Admix [18], SSA [19], STM [72], USMM [73], DeCoWA [74], L2T [20], BSR [75] and OPS [21]); (2) 8 model-driven methods comprising model ensemble (SVRE [24], CWA [25]) and surrogate refinement (SGM [22], LinBP [23], AWT [76], FAUG [77], ANA [78], LL2S [79]) strategies; (3) 24 optimization-driven methods including feature-based (ILA [26], NAA [27], BFA [80], P2FA [81]), generative (CDA [28], DiffAttack [82]), gradient stabilization (MI [16], NI [17], PI/EMI [4], VMI/VNI [1], ANDA [44], FGSRA [45], GI [46], MUMODIG [47], GAA [48], FoolMix [49] ), and flatness-enhanced (RAP [7], PGN [8], GNP [5], FEM [3], APP [6], TPA [2]) approaches. PGN [8] constitutes the state-of-the-art in optimization-driven methods. We further benchmark against 4 classical query-based attacks (SimBA [13], Square [83], Boundary [12] and RayS [84]) to provide a holistic view under different threat models.
Hyper-parameters. All baseline methods use their originally reported configurations unless specified (see Supp. -C). We unify common parameters across attacks: maximum perturbation $\epsilon=16/255$, and step size $\alpha=\epsilon/T$. Our MEF sets neighborhood radius $\gamma=2\times\epsilon$, exploration radius $\xi=0.15\times\epsilon$, outer/inner momentum coefficient $\mu_{outer}/\mu_{inner}=0.5/0.9$. MEF requires only one backpropagation per sample versus PGN (current SOTA)’s two. For comprehensive evaluation, we design two variants: MEFH (10 iterations, half PGN’s cost) and MEFF (20 iterations, fully matching PGN’s cost).
VI-B Effectiveness Benchmarking against Existing Attacks
VI-B1 Comparison with Gradient Stabilization Attacks
We benchmark MEF against 18 representative gradient stabilization attacks on normally trained models. We generate adversarial examples on four source models and evaluate their transferability across eight diverse target models.
Table II demonstrates the significant superiority of our framework. To maintain clarity and focus on the state-of-the-art, Table II reports methods with an average success rate exceeding 80%. Results for other classical baselines (e.g., MI [16], RAP [7]) are provided in Supp. -D (Table A-1). We compute the overall average success rate by averaging the AVG columns across all four surrogate settings. Although FEM [3] shows competitive performance on ResNet-based surrogates, PGN remains the second-best method overall with a higher global average (88.20% vs. 88.13% for FEM [3]). Notably, MEFH attains 89.93% average success, surpassing the state-of-the-art PGN [8] (88.20%) despite requiring only half the budget. Matching PGN [8]’s cost, MEFF elevates the average success rate to 94.84%, establishing MEF as the new state-of-the-art for adversarial transferability.
TABLE III: Transfer attack success rate ($\%$) comparison on diverse architectures. best results are bold and second best are underlined.
Attack Res-50 $\Longrightarrow$ Inc-v3 $\Longrightarrow$
MobileNet PNASNet-L ViT-B/16 ViT-L/32 PiT-S MLP-Mixer ResMLP MobileNet PNASNet-L ViT-B/16 ViT-L/32 PiT-S MLP-Mixer ResMLP
FEM [3] 94.0 87.0 43.1 35.0 58.1 60.2 66.0 81.1 78.2 41.3 30.4 53.2 59.7 58.6
PGN [8] 91.6 88.6 51.4 45.8 58.7 66.7 70.2 84.9 82.6 50.1 40.6 60.2 66.1 65.6
FGSRA [45] 88.8 87.0 49.8 45.6 57.1 64.5 70.9 83.7 81.7 48.1 39.6 59.6 64.2 65.2
MUMODIG [47] 89.5 86.3 48.8 44.1 56.9 66.3 69.0 81.3 81.1 48.6 39.0 59.5 65.7 65.1
MEFH 95.2 90.3 51.9 46.0 63.3 66.9 70.5 85.2 83.9 51.7 42.2 60.4 67.3 66.3
MEFF 98.0 95.2 56.0 46.9 71.1 71.7 77.4 92.7 90.8 56.2 45.3 71.4 69.4 73.3
VI-B2 Evaluation on Diverse Network Architectures
To assess cross-architecture transferability, we evaluate attack success rates on seven heterogeneous models, including MobileNet [58], PNASNet [59], Vision Transformers (ViT) [60, 61], and MLP-Mixers [62, 63]. Table III compares MEF against the top-4 performing baselines (full comparison with 12 baselines is in Supp. -G). Despite the significant inductive bias gap, MEFF consistently outperforms the state-of-the-art PGN [8], achieving a remarkable 46.9% success rate on ViT-L/32 (vs. PGN’s 45.8%) and 77.4% on ResMLP (vs. PGN’s 70.2%). This confirms that optimizing zeroth-order flatness effectively extracts geometry-invariant features that transfer across fundamentally different architectural paradigms.
TABLE IV: Transfer attack success rate (%) comparison on ensemble surrogates. Best results are bold and second best are underlined.
Attack Res-50 + Inc-v3 $\Longrightarrow$
Res-101 Inc-v4 IncRes-v2 Inc-v3adv Inc-v3ens3 Inc-v3ens4 IncRes-v2ens MobileNet PNASNet-L ViT-B/16 ViT-L/32 PiT-S MLP-Mixer ResMLP
FEM [3] 98.7 92.0 88.3 68.3 65.9 62.9 46.9 92.5 89.0 59.0 41.3 74.1 66.0 77.0
PGN [8] 97.8 91.9 89.4 79.5 78.2 77.5 65.6 93.9 91.9 68.7 50.2 76.9 74.5 81.8
FGSRA [45] 94.9 90.3 87.4 78.9 76.4 75.2 66.3 92.6 90.9 66.5 49.1 76.1 72.5 81.2
MUMODIG [47] 95.5 89.6 86.0 77.0 76.0 76.9 64.5 90.1 90.2 66.8 48.4 75.8 73.9 80.9
MEFH 98.9 93.2 90.2 80.2 79.8 80.5 67.0 94.1 92.4 69.5 51.7 77.8 75.3 82.1
MEFF 99.4 96.6 95.1 85.7 83.3 82.6 70.4 96.7 95.1 75.4 55.2 83.8 80.4 89.0
TABLE V: Transfer attack success rate (%) comparison with input augmentation attacks. The best results are bold.
Attack Res-50 $\Longrightarrow$ Inc-v3 $\Longrightarrow$
IncRes-v2 Inc-v3ens3 Inc-v4ens4 IncRes-v2ens ViT-L/32 MLP-Mixer AVG IncRes-v2 Inc-v3ens3 Inc-v4ens4 IncRes-v2ens ViT-L/32 MLP-Mixer AVG
TI [15] 51.1 42.0 39.1 30.2 34.4 55.6 42.1 52.0 37.0 37.2 26.0 28.8 53.7 39.1
MEFH+TI 81.7 77.4 77.8 68.2 57.3 71.1 72.2 86.4 79.6 79.5 67.4 48.7 68.3 71.6
MEFF+TI 94.1 83.5 82.6 73.6 62.5 76.0 78.7 94.5 89.0 88.3 78.8 56.3 75.6 80.4
STM [72] 84.9 52.0 49.1 32.9 37.5 66.4 53.8 85.6 54.2 56.4 32.7 34.9 63.8 54.6
USMM [73] 84.8 51.7 48.8 32.6 36.9 66.0 53.5 85.4 53.4 56.3 32.3 34.9 63.3 54.3
DeCowA [74] 90.1 59.4 56.7 40.6 44.3 70.2 60.2 92.9 62.3 63.9 40.2 42.8 71.6 62.3
L2T [20] 89.6 56.8 53.8 37.5 41.6 69.7 58.2 90.3 58.5 61.1 37.0 40.1 69.3 59.4
BSR [75] 89.4 57.9 55.0 38.6 43.1 69.9 59.0 91.4 59.5 62.2 38.2 40.1 70.0 60.2
OPS(10,5,5) [21] 96.2 84.3 81.9 72.3 63.7 77.7 79.3 95.9 90.7 88.8 79.4 57.1 76.3 81.4
OPS(10,30,30) [21] 98.5 90.6 89.2 83.7 81.1 98.1 90.2 98.7 93.1 93.2 86.6 75.3 91.9 89.8
MEFH 83.7 64.5 63.3 50.4 40.8 65.3 61.3 86.1 65.1 65.4 45.4 35.2 65.5 60.4
MEFF 90.3 69.2 65.1 52.2 44.9 71.7 65.6 94.3 69.0 69.6 47.9 40.3 69.4 65.1
[…]
VI-B3 Attacking on Ensemble of Models
We extend our evaluation to ensemble attacks by integrating Inc-v3 [37] and Res-50 [9] through logit averaging as the surrogate model. Table IV compares MEF against the top-4 performing baselines (full comparison with 14 methods, including ensemble-specialized SVRE [24] and CWA [25], is provided in Supp. -H). MEFF demonstrates consistent superiority, outperforming the state-of-the-art PGN [8] by margins of 1.6%-8.2% across diverse targets. This confirms that MEF captures transferable adversarial patterns across model ensembles.
VI-B4 Integrated with Input Augmentation Attacks
We compare MEFH and MEFF against 10 representative input augmentation attacks (e.g., SSA [19], DeCoWA [74]). Following the standard setup, we generate adversarial examples on Res-50 [9] and Inc-v3 [37] and evaluate their transferability to challenging targets, including ensemble models and heterogeneous architectures. Furthermore, to demonstrate compatibility, we integrate sample transformations from five classical augmentation methods (e.g., TI [15], DI [14]) into MEF variants without introducing additional gradient computations.
[…]
TABLE VI: Transfer attack success rate (%) comparison across attack paradigms. The best results are bold.
Attack Inc-v3 Inc-v4 Res-101 Inc-v3ens3 Inc-v3ens4 IncRes-v2ens AVG
SGM [22] 71.6 35.3 92.8 12.3 12.5 6.1 38.4
LinBP [23] 72.7 47.9 95.4 13.2 13.3 6.8 41.6
AWT [76] 89.1 87.5 97.9 63.1 61.4 49.1 74.7
FAUG [77] 68.0 64.1 78.6 41.5 42.2 27.4 53.6
ANA [78] 79.8 75.4 86.0 51.4 50.3 38.9 63.6
LL2S [79] 88.2 86.6 96.7 61.9 60.6 47.8 73.6
ILA [26] 70.5 74.8 97.9 16.5 15.3 9.0 47.3
NAA [27] 82.7 82.4 98.3 30.1 29.6 21.0 57.4
BFA [80] 89.1 87.3 98.6 63.2 60.9 48.6 74.6
P2FA(10,70) [81] 95.4 95.7 99.8 70.1 66.4 54.2 81.5
CDA [28] 69.4 83.1 94.5 49.8 52.7 43.3 65.5
DiffAttack [82] 41.4 41.9 52.6 15.9 14.0 5.4 28.5
MEFH 90.6 88.7 99.7 64.5 63.3 50.4 76.2
MEFF 95.1 94.2 99.9 69.2 65.1 52.2 79.3
MEF(10,70) 97.5 96.6 99.9 73.1 68.5 55.7 83.6
VI-B5 Comparison with Other Types of Transfer-based Attacks
We validate MEF against three paradigms: surrogate refinement (SGM [22], LinBP [23], AWT [76], FAUG [77], ANA [78], LL2S [79]), feature-based (ILA [26], NAA [27], BFA [80], P2FA [81]), and generative attacks (CDA [28], DiffAttack [82]). As shown in Table VI, MEF consistently outperforms these approaches on Res-50 [9]. Notably, MEFF (79.3%) surpasses leading baselines like AWT [76] (74.7%) and BFA [80] (74.6%) solely through geometry-aware input space optimization, avoiding reliance on architectural priors.
Among feature-based methods, the state-of-the-art method P2FA [81] attains 81.5% but incurs $3.5\times$ higher computational cost. Under a budget-aligned setting, our MEF(10,70) variant reclaims the lead with 83.6% (+2.1% over P2FA), confirming the fundamental superiority of zeroth-order flatness optimization over intermediate feature manipulation.
TABLE VII: Comparison of Efficiency, Constraint Compliance, and Stealthiness on Inc-v3 ($\epsilon=16/255$).
Method Category Queries ASR (%) LPIPS $\downarrow$ SSIM $\uparrow$
Boundary [12] Decision-based 500 12.3 0.85 0.291
Boundary [12] Decision-based 5,000 24.5 0.72 0.346
RayS [84] Decision-based 500 59.8 0.18 0.887
RayS [84] Decision-based 5,000 94.3 0.05 0.947
SimBA [13] Score-based 500 7.80 0.01 0.991
SimBA [13] Score-based 5,000 41.8 0.01 0.992
Square [83] Score-based 500 94.9 0.31 0.583
Square [83] Score-based 5,000 99.7 0.30 0.579
MEFH Transfer-based 0 90.6 0.24 0.751
MEFF Transfer-based 0 95.1 0.30 0.692
VI-B6 Comparison with Query-based Attacks
We benchmark MEF against score-based (SimBA [13], Square [83]) and decision-based (Boundary [12], RayS [84]) attacks on Inc-v3 [37]. Unlike feedback-dependent baselines, MEF operates in a zero-query setting using a ResNet-50 surrogate. As shown in Table VII, while Square Attack [83] achieves marginally higher success (99.7%), it incurs prohibitive costs (5,000 queries) and severe visual degradation (SSIM 0.579). In contrast, our zero-query $\text{MEF}_{\text{F}}$ attains a competitive 95.1%, outperforming the decision-based SOTA RayS [84] (94.3% at 5,000 queries). Even the half-cost $\text{MEF}_{\text{H}}$ maintains 90.6% success with superior stealth (SSIM 0.751). This confirms MEF offers a practical solution without the detection risks and computational overhead of high-frequency queries.
VI-C Practical Threats to Real-World Applications
TABLE VIII: Transfer Attack Success Rates on Six Commercial Vision APIs. The best results are bold and the second best are underlined.
Method Google Cloud Vision Amazon Rekognition Azure AI Vision
T-1 T-5 T-10 T-1 T-5 T-10 T-1 T-5 T-10
PGN [8] 67.20% 44.30% 40.00% 54.50% 36.60% 26.30% 59.10% 32.00% 27.40%
FEM [3] 65.30% 44.10% 39.00% 57.10% 40.00% 31.00% 54.40% 26.10% 21.20%
FGSRA [45] 68.60% 46.10% 40.90% 55.90% 39.70% 29.20% 51.70% 22.30% 18.50%
MEFH 71.60% 50.20% 46.10% 58.50% 42.80% 33.90% 60.90% 34.70% 29.10%
MEFF 79.40% 63.30% 59.00% 68.80% 55.50% 46.70% 67.18% 40.33% 34.88%
Method Baidu AI Cloud Alibaba Cloud Tencent Cloud
T-1 T-5 T-10 T-1 T-5 T-10 T-1 T-5 T-10
PGN [8] 79.80% 61.10% 61.10% 86.40% 66.10% 66.10% 74.90% 53.10% 51.90%
FEM [3] 80.60% 61.30% 61.30% 83.40% 64.20% 64.20% 72.50% 50.80% 49.80%
FGSRA [45] 74.50% 57.80% 57.80% 83.00% 66.00% 66.00% 76.00% 52.70% 51.50%
MEFH 83.40% 62.70% 62.70% 86.80% 66.80% 66.80% 75.10% 54.70% 53.60%
MEFF 86.30% 70.10% 70.10% 91.70% 77.30% 77.30% 84.70% 67.10% 66.10%
VI-C1 Evaluation on Commercial Vision APIs
To assess practical threats beyond open-source models, we evaluate transferability against six commercial black-box systems. Unlike local models, these APIs run proprietary architectures with unknown preprocessing defenses. Given query costs, we compare MEF against the three leading baselines using the Top-k Exclusion (T-k) metric, where the correct ground-truth label is excluded from the top-k predictions.
Table VIII demonstrates that MEF consistently outperforms baselines across all platforms. Notably, on the rigorous Google Cloud Vision API, MEFF achieves 79.40% T-1, surpassing the SOTA PGN (67.20%) by a significant margin of 12.2%. Even the half-cost MEFH (71.60%) exceeds the strongest baseline. High success rates are also observed on Alibaba Cloud (91.70%) and others. These results confirm that the flat local minima identified by our framework generalize effectively to unknown production pipelines. For intuitive visual evidence of these successful semantic subversions, we provide qualitative comparisons of API outputs in Fig. A-4 of the Supp. -L.
TABLE IX: Transferability comparison (Fooling Rate %) against Multimodal Large Models. The best results are bold and the second best are underlined
Attack Method LLaVA Unidiffuser MiniGPT-4 Qwen2.5-VL Average
APP [6] 19.1 24.9 23.8 23.4 22.8
FEM [3] 22.5 29.3 26.5 29.4 26.9
PGN [8] 22.0 32.4 27.8 24.5 26.7
FGSRA [45] 20.3 26.4 25.7 24.7 24.3
ANDA [44] 19.6 25.5 24.9 23.9 23.5
GI [46] 19.8 25.7 25.0 24.1 23.6
MUMODIG [47] 20.2 26.2 25.5 24.5 24.1
GAA [48] 19.8 25.8 25.1 24.1 23.7
FoolMix [49] 19.4 25.2 24.5 23.6 23.2
MEFH 26.7 35.4 30.5 33.4 31.5
MEFH 30.5 38.4 35.5 41.5 36.5
VI-C2 Transferability to Multimodal Large Language Models
To investigate transferability against Multimodal Large Language Models (MLLMs), we evaluate MEF on LLaVA [68], Unidiffuser [69], MiniGPT-4 [70], and Qwen2.5-VL [71]. We generate adversarial examples on Res-50 [9] and transfer them to these models. This challenging cross-task scenario requires disrupting semantic alignment across modalities.
We prompt models with “Describe the image” to generate captions $C_{cln}$ and $C_{adv}$ for clean and adversarial images. Let $D(\cdot)$ denote the CLIP text encoder [85]. We report the CLIP-based Fooling Rate, which is defined as the proportion of samples satisfying the semantic divergence condition:
$\text{CosSim}\big(D(C_{cln}),D(C_{adv})\big)\leq\tau,$ (15)
where $\tau=0.4$ is an empirical threshold adopted to distinguish successful attacks from benign caption variations.
As shown in Table IX, MEF demonstrates superior cross-paradigm transferability. MEFF achieves the highest average fooling rate of 36.5% across all models. Notably, even the efficient MEFH (31.5% average) outperforms the best baseline FEM (26.9%). On Qwen2.5-VL, MEFF reaches 41.5%, surpassing baselines by over 12%. This confirms that optimizing multi-order flatness creates robust adversarial features capable of compromising the visual encoders of advanced MLLMs (see Supp. -I for qualitative examples).
VI-D Robustness against Defense Mechanisms
TABLE X: Transfer attack success rate ($\%$) against modern defense mechanisms. Evaluating across three defense categories on seven protected targets, our MEFH/F outperform all competitors. The best results are bold and the second best are underlined.
Attack Res-50 $\Longrightarrow$ Inc-v3 $\Longrightarrow$
Inc-v3adv Inc-v3ens3 Inc-v3ens4 IncRes-v2ens RS HGD NRP AVG Inc-v3adv Inc-v3ens3 Inc-v3ens4 IncRes-v2ens RS HGD NRP AVG
APP [6] 50.1 49.0 48.5 34.6 38.6 60.3 88.0 52.7 52.7 46.2 46.3 28.9 30.4 23.1 43.2 38.7
FEM [3] 50.2 49.3 46.4 31.4 36.4 72.2 89.5 53.6 56.2 49.7 49.4 28.5 29.9 23.5 43.0 40.0
PGN [8] 60.2 58.4 59.6 45.6 48.1 62.5 92.9 61.0 69.2 61.1 62.2 41.6 38.1 27.7 54.6 50.6
ANDA [44] 50.2 48.9 47.3 37.7 41.0 62.5 87.7 53.6 58.2 50.2 50.7 29.5 29.4 22.3 45.1 40.8
FGSRA [45] 52.8 51.4 52.1 36.7 39.8 63.6 91.6 55.4 57.9 50.5 51.1 31.9 30.5 26.1 45.2 41.9
GI [46] 50.5 50.9 48.1 37.1 37.9 60.7 89.7 53.6 57.5 51.3 52.2 32.0 32.1 22.5 44.1 41.7
MUMODIG [47] 53.2 51.6 51.8 36.9 40.2 63.7 87.7 55.0 57.2 51.0 52.1 32.2 32.1 23.0 46.1 42.0
GAA [48] 49.7 49.0 51.1 36.9 39.5 62.8 85.6 53.5 55.2 49.7 50.0 31.5 31.7 25.0 45.1 41.2
FoolMix [49] 49.6 50.7 46.2 34.9 38.1 62.2 84.5 52.3 55.6 48.5 49.5 29.7 32.3 23.4 42.3 40.2
MEFH 62.4 64.5 63.3 50.4 49.2 80.6 94.6 66.4 69.7 65.1 65.4 45.4 38.3 46.3 54.8 55.0
MEFF 70.1 69.2 65.1 52.2 51.5 83.0 97.2 69.8 75.8 69.0 69.6 47.9 40.4 52.4 56.2 58.8
VI-D1 Evaluation on Standard Defense Methods
To evaluate robustness, we transfer adversarial examples generated on Inc-v3 and Res-50 to seven representative protected targets spanning adversarial training [29, 30], purification [32, 33], and certifiable defense [31], as detailed in Sec. IV-A. Table X compares MEFH/MEFF with competitive gradient stabilization attacks that achieve an average success rate exceeding 40%. Results for other baselines (e.g., MI [16], VMI [1]) are provided in Supp. -F (Table A-3). MEFH achieves 60.70% average success rate across all defenses using half the computational budget of PGN [8], a 4.86% improvement over PGN’s 55.84% baseline. When matching PGN’s cost, MEFF further elevates the average success rate to 64.30%, a 8.46% absolute improvement. Notably, MEFF outperforms all competitors on individual defense types. This consistent superiority demonstrates our method’s capability to bypass diverse defense mechanisms, establishing new state-of-the-art robustness against protected models.
TABLE XI: Transfer attack success rate ($\%$) against advanced defenses. The best results are bold and the second best are underlined.
Attack Diffusion-based Purifier RobustBench Models
DiffPure[34] DensePure[35] MIMIR[65] ARES[66] MeanSparse[67]
 Surrogate Model: ResNet-50
APP [6] 40.2 40.0 8.1 9.1 9.9
FEM [3] 42.8 42.0 8.1 9.1 9.9
PGN [8] 46.9 46.7 9.5 10.7 11.6
ANDA [44] 39.3 39.1 8.0 8.9 9.7
FGSRA [45] 41.3 43.7 9.1 10.2 11.1
GI [46] 40.9 40.7 8.3 9.3 10.1
MUMODIG [47] 41.4 41.3 8.4 9.4 10.2
GAA [48] 39.4 39.2 8.0 9.0 9.7
FoolMix [49] 40.7 40.5 8.3 9.3 10.1
MEFH 44.5 44.3 8.8 9.8 10.7
MEFF 48.6 47.2 17.3 19.4 21.1
 Surrogate Model: Inception-v3
APP [6] 34.8 34.4 7.0 7.9 8.6
FEM [3] 41.6 40.9 8.4 9.4 10.2
PGN [8] 45.1 44.5 9.1 10.2 11.1
ANDA [44] 37.1 36.6 7.5 8.4 9.1
FGSRA [45] 44.9 44.2 8.7 9.8 10.6
GI [46] 37.9 37.4 7.6 8.6 9.3
MUMODIG [47] 37.6 37.1 7.6 8.5 9.3
GAA [48] 36.7 36.2 7.4 8.3 9.0
FoolMix [49] 35.8 35.3 7.2 8.1 8.8
MEFH 43.7 42.9 8.5 9.6 10.4
MEFF 46.1 45.1 15.8 17.8 19.3
VI-D2 Evaluation on Advanced State-of-the-Art Defenses
To address the rapid evolution of defenses, we extend evaluation to state-of-the-art mechanisms: diffusion-based purification (DiffPure [34], DensePure [35]) and top-performing robust models from RobustBench [64] (MIMIR [65], ARES [66], MeanSparse [67]). For comparison, we select the top-9 performing baselines from Table X. As shown in Table XI, these advanced defenses pose significant challenges. Nevertheless, MEFF consistently demonstrates superior resilience. Against purification defenses, it outperforms the runner-up PGN [8] by 1.7% on DiffPure [34]. More notably, against the highly robust MeanSparse model [67], MEFF achieves a remarkable success rate of 21.1% (ResNet-50 [9] as surrogate model), nearly doubling PGN [8]’s 11.6%. This suggests that MEF’s average-case flatness optimization generates adversarial examples that are structurally stable enough to survive stochastic purification and robust feature alignment.
VI-E Analysis of Efficiency, Stealthiness, and Sensitivity
VI-E1 Convergence Analysis and Iteration Selection
Figure 5: Attack Success Rate (ASR) convergence curve over 100 iterations (Source: Res-50 [9] $\rightarrow$ Target: Inc-v3 [37]). The dashed red line marks the conventional $T=10$ iteration cutoff.
TABLE XII: Convergence and Performance Comparison. $\Delta$ denotes the improvement from $T_{10}$ to $T_{max}$. The best results are bold.
Method $\mathbf{T_{10}}$ ASR (%) $\mathbf{T_{\text{max}}}$ ASR (%) $\mathbf{\Delta}$ ASR (%) $\mathbf{T_{\text{max}}}$ Iter.
MI [16] 50.50% 51.20% 0.70% 56
NI [17] 55.70% 56.60% 0.90% 12
PI [4] 61.00% 61.00% 0.00% 10
FEM [3] 89.50% 89.60% 0.10% 12
FGSRA [45] 87.10% 89.70% 2.60% 23
PGN [8] 86.00% 88.50% 2.50% 40
MEF (Ours) 89.80% 96.40% 6.60% 36
To evaluate the impact of iteration count $T$ on transferability, we conducted an extended convergence analysis ($T=100$) using Res-50 [9] as the surrogate and Inc-v3 [37] as the target. We selected six representative baselines, comprising both classic and SOTA methods. Since attackers in realistic black-box scenarios lack target model feedback to identify the optimal stopping point, assessing whether attacks can sustain transferability over extended iterations is crucial.
The results, visualized in Fig. 5 and detailed in Table XII, reveal distinct behaviors. For most baselines, including the SOTA PGN [8], performance rapidly plateaus or even declines after reaching an early peak. This trend aligns with established findings [16], indicating that excessive iterations lead to overfitting on the surrogate model, thereby hampering transferability. Consequently, the gap between their $T_{10}$ performance and absolute peak ($T_{max}$) is marginal ($3\%$), validating $T=10$ as a cost-effective standard for these methods. In contrast, MEF effectively mitigates overfitting through flatness optimization, exhibiting a significantly higher optimization ceiling. It continues to accumulate gains up to $T=36$, achieving a peak ASR of 96.40% (a 6.6% improvement over $T=10$). Crucially, MEF remains strictly superior to all competitors throughout the entire 100-iteration process, confirming its dominance for both resource-constrained ($T=10$) and determined attackers.
VI-E2 Perceptual Quality Analysis
Figure 6: Comparison of Transferability (ASR) vs. Visual Quality (SSIM). Transfer ASR are valuated on Res-50 [9] $\to$ Inc-v3 [37] with $\epsilon=16/255$.
TABLE XIII: Quantitative comparison of visual perceptibility. The best results are bold and the second best are underlined.
Method Transferability Visual Quality Metrics
ASR (%) $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$
MUMODIG [47] 84.4 0.696 26.86 0.323 0.247
FGSRA [45] 85.3 0.683 26.62 0.313 0.252
PGN [8] 86.3 0.681 26.36 0.294 0.250
FEM [3] 89.4 0.666 26.44 0.322 0.251
MEFH 90.6 0.721 28.35 0.250 0.241
MEFF 95.1 0.692 26.56 0.295 0.244
We evaluate perceptual quality ($\epsilon=16/255$) via the ASR-SSIM trade-off (Fig. 6) and quantitative metrics (Table XIII). As shown in Fig. 6, MEF occupies the optimal top-right quadrant, balancing strength and stealth. Quantitatively, $\text{MEF}_{\text{H}}$ achieves the best fidelity (SSIM 0.721, LPIPS 0.250), surpassing the prior leader MUMODIG [47]. Meanwhile, $\text{MEF}_{\text{F}}$ yields the highest transferability (95.1%) with competitive stealth (outperforming PGN [8] in DISTS). This confirms MEF’s superiority stems from geometric efficiency. Qualitative results are in Supp. -J.
VI-E3 Computational Efficiency Analysis
TABLE XIV: Computational efficiency and resource cost comparison. We report computational cost (Backpropagations/BP), Peak GPU Memory (MiB), and average runtime (seconds/image).
Method Time (s/img) $\downarrow$ BP (count) $\downarrow$ Peak Mem. (MiB) $\downarrow$ ASR (%) $\uparrow$
MI [16] 0.29 10 1397 50.5
NI [17] 0.29 10 1400 56.5
PI [4] 0.30 10 1403 60.2
VMI [1] 0.56 200 4316 72.8
VNI [1] 0.59 200 4315 77.2
EMI [4] 0.48 200 4309 75.0
FEM [3] 2.80 210 1,788 89.4
FGSRA [45] 0.96 400 4100 85.3
PGN [8] 1.13 400 4488 86.3
MEFH 0.46 200 4090 90.6
MEFF 0.97 400 4097 95.1
We evaluate computational efficiency by measuring the hardware-agnostic Backpropagation (BP) counts, Peak GPU Memory, and the average wall-clock runtime required to generate an adversarial example (averaged over 100 randomly selected samples on an NVIDIA RTX 4090). For fair comparison, all sampling-based methods employ $N=20$ points with full synchronous parallel computation. We benchmark against basic single-point attacks and representative, well-established sampling-based methods.
As detailed in Table XIV, single-point methods are naturally the fastest (10 BPs, $\approx$ 1400 MiB) but suffer from limited transferability. In contrast, MEFH achieves a remarkable 90.6% success rate with only 200 BPs and 4090 MiB memory. Crucially, compared to the flatness-based SOTA PGN [8] (400 BPs, 4488 MiB), MEFH halves the gradient cost and reduces memory footprint, delivering a $\mathbf{2.4\times}$ speedup while improving ASR by 4.3%. Even our full-cost variant MEFF remains faster than PGN with lower memory usage, and significantly outpaces heavy methods like FEM [3] (2.80s). This confirms that our GBO strategy effectively mitigates the computational burden, achieving an optimal efficiency-performance ratio.
(a) $\gamma$
(b) $\xi$
(c) NN
(d) $\mu_{outer}$
(e) $\mu_{inner}$
(f) $T$
Figure 7: Parameter ablation study on MEF, evaluating $\gamma$$\xi$, $\mu_{inner}$$\mu_{outer}$, and NN/$T$ impacts on transfer success rates across six challenging targets.
VI-E4 Sensitivity to Perturbation Budget
To assess attack robustness under stricter imperceptibility constraints, we evaluate MEF against the top-4 performing baselines (Res-50 $\to$ Inc-v3) while varying $\epsilon$ from $2/255$ to $16/255$. As illustrated in Fig. 8, MEF consistently outperforms competitors across all magnitudes. Notably, at the restrictive budget of $\epsilon=4/255$, MEF achieves a success rate of 41.3%, surpassing the runner-up PGN (31.7%) by a significant margin of 9.6%. This confirms that MEF’s superiority stems from the precise identification of geometrically flat directions rather than reliance on large perturbation norms. For detailed visual inspections across these budgets, please refer to Supp. -J.
Figure 8: Attack success rate under varying perturbation budgets ($\epsilon$).
VI-E5 Parameter Sensitivity
To analyze the impact of key components in MEF, we evaluate six parameters ($\gamma,\xi,\mu_{outer},\mu_{inner},N,T$) across seven diverse targets, ranging from standard CNNs to defended models. The results illustrated in Figure 7 reveals three trends. First, both neighborhood radius $\gamma=2\epsilon$ and exploration radius $\xi=0.15\epsilon$ generalize as optimal values across all targets, demonstrating that MEF is robust to parameter choices and does not require laborious model-specific tuning. Second, inner momentum $\mu_{inner}=0.9$ proves indispensable (delivering a 10% boost), while outer momentum provides marginal gains. Third, increasing sampling size $N$ shows continuous improvement, while extending iterations $T$ to 20 yields substantial gains (37.2%).
TABLE XV: Quantitative Comparison of Zeroth/First-Order Flatness.
Attack Res-50 [9] Inc-v3 [37]
($\overline{R}^{(0)}/\overline{R}^{(1)}$) ($\overline{R}^{(0)}/\overline{R}^{(1)}$)
VMI [1] 29.03 / 0.49 8.32 / 0.64
VNI [1] 23.04 / 0.41 7.74 / 0.59
EMI [4] 25.57 / 0.45 8.77 / 0.62
RAP [7] 15.98 / 0.45 7.80 / 0.62
GNP [5] 50.49 / 0.53 29.37 / 0.86
FEM [3] 12.20 / 0.55 3.52 / 0.79
APP [6] 2.41 / 0.47 1.75 / 0.65
TPA [2] 2.71 / 0.18 1.24 / 0.27
PGN [8] 2.93 / 0.37 2.11 / 0.56
MEFH 2.20 / 0.26 1.09 / 0.27
MEFF 2.34 / 0.10 1.06 / 0.21
VI-E6 Landscape Flatness Analysis
[…]
VII Discussion and Conclusion
Discussion. While MEF significantly outperforms state-of-the-art baselines across challenging cross-task and cross-architecture settings, strictly bridging the structural gap between surrogate and target models remains an open challenge. Moreover, we acknowledge the inherent trade-off between transferability and visual stealthiness in the zero-query setting. Unlike white-box or query-based approaches, transfer-based attacks generally necessitate a larger perturbation budget to ensure the generalization of adversarial features to unknown targets. Nevertheless, MEF establishes a superior trade-off, achieving the most favorable balance between ASR and visual quality among existing transfer-based methods. Addressing these limitations constitutes a meaningful direction for future research. Specifically, future exploration of agnostic priors aims to reconcile the trade-off between cross-domain transferability and visual stealthiness.
Conclusion. This work establishes the first theoretical foundation linking multi-order flatness to adversarial transferability. By unifying fragmented definitions and resolving the exploitation-exploration dilemma, we proposed the Maximin Expected Flatness (MEF) attack. Extensive evaluations demonstrate that MEF achieves state-of-the-art transferability and superior efficiency across diverse models and defenses. We hope this work inspires both deeper and broader adoption of principled flatness optimization in AI security.